# A Lot of Talk and a Badge: An Exploratory Analysis of Personal Achievements in GitHub

### 30 Dec 2023 v2

@@ Abstract
Context. GitHub has introduced a new gamification element through personal achievements, whereby badges are unlocked and displayed on developers’ personal profile pages in recognition of their development activities. Objective. In this paper, we present an exploratory analysis using mixed methods to study the diffusion of personal badges in GitHub, in addition to the effects and reactions to their introduction. Method. First, we conduct an observational study by mining longitudinal data from more than 6,000 developers and performed correlation and regression analysis. Then, we conduct a survey and analyze over 300 GitHub community discussions on the topic of personal badges to gauge how the community responded to the introduction of the new feature. Results. We find that most of the developers sampled own at least a badge, but we also observe an increasing number of users who choose to keep their profile private and opt out of displaying badges. Besides, badges are generally poorly correlated with developers’ qualities and dispositions such as timeliness and desire to collaborate. We also find that, except for the Starstruck badge (reflecting the number of followers), their introduction does not have an effect. Finally, the reaction of the community has been in general mixed, as developers find them appealing in principle but without a clear purpose and hardly reflecting their abilities in the current form. Conclusions. We provide recommendations to GitHub platform designers on how to improve the current implementation of personal badges as both a gamification mechanism and as sources of reliable cues of ability for developers’ assessment.
@@ 1 Introduction
Gamification refers to the practice of incorporating elements and mechanics of game design into non-game contexts (Hamari, 2007; Hunter and Werbach, 2012). The goal of gamification is to increase user engagement and motivation, as well as drive desired behaviors by creating a sense of play and/or competition (Deterding et al., 2011; Robson et al., 2015). There have been several attempts to propose frameworks for the gamification of Software Engineering activities (García et al., 2017; Sasso et al., 2017) and methods to build gamified software (Morschheuser et al., 2018). Previous research has also found evidence that gamification mechanisms are in part responsible for the success of websites such as Stack Overflow (Vasilescu et al., 2013). Gamification, in fact, has always been a staple of the technical Q&A site whose members gain badges, medals, and privileges by being good community citizens (e.g., answering questions, editing poorly written posts, etc.). Another popular source of user-generated content in the Software Engineering domain is GitHub, which, in contrast, has always incorporated several social features (e.g., developers have their personal profiles and can follow each other, repositories can be starred), but only implemented few and simple gamification elements on its website, such as code streaks counters, which were removed from the contribution graph in May 2016 (Vessels, 2016). However, this suddenly changed in June 2022, when a blog post (Brooks, 2022) announced the introduction of personal achievements, a new feature whereby badges are unlocked in recognition of developers’ activities and displayed on their personal profile pages. However, unlike Stack Overflow, neither the list of available badges is publicly disclosed by GitHub nor their meaning and the rules to unlock them. Previous research has shown that implementing gamification elements in collaborative software platform may steer the behavior of developers in unexpected and unwanted directions (Moldon et al., 2021). In this work, We find evidence that it is not entirely clear to the GitHub community how to interpret the meaning of personal achievements. As shown in Figure 1, a user is asking the community for help (1) on how to interpret the Pull Shark badge (2), and two other users agree with the original poster that the meaning of the badge is ambiguous and not intuitive (3).
@@
GitHub has not only earned the reputation of being the one-stop-shop for software development (Metz, 2015) but has also quickly become a useful platform for job seekers (sup, 2022; hac, 2011; ant, 2012). Employers access applicants’ GitHub profile pages to supplement resumes and further assess the knowledge, skills, and abilities of candidates (Marlow and Dabbish, 2013). Marlow and Dabbish (2013) found that GitHub design strongly influenced the activity traces available through profile pages, which employers look at during applicants’ assessment. In particular, they observed that employers look at those activity traces that provide reliable cues and require lower effort to access and evaluate them. Because personal achievement badges are gamification elements that provide quick and low-effort access to activity traces, GitHub users who display them on their profile pages may be unintentionally ‘giving off’ cues that have unexpected repercussions, for example, to managers assessing them as job candidates and to other OSS developers who review their contributions (Singer et al., 2013; Tsay et al., 2014). In this study, we find evidence that few of the activity cues conveyed through personal badges are actually associated with the intended developers’ abilities.
@@
In this paper, we investigate the effects of introducing a new gamification element into GitHub. Specifically, we conduct a mixed-methods study of personal achievement badges (in short, personal badges hereinafter). We characterize the distribution of badges in GitHub (RQ), investigate the qualities of developers they are intended to signal and how well they correlate with these qualities (RQ), and look into the effects, if any, that the introduction of this feature has had on developers’ activities (RQ). Finally, we gauge how the GitHub community feels about the new feature (RQ).
@@
After mining longitudinal data from over 6,000 GitHub developers, we find that while most own at least one badge, the number of those who chose to have a private page or opted out of displaying badges on their profile has been steadily increasing. Furthermore, the correlation and regression analyses reveal that, respectively, only one out of the five types of personal badges analyzed (i.e., Starstruck) provides a reliable signal of the associated developer quality (i.e., popularity) and that its introduction has positively affected the number of followers for developers displaying the badge. Finally, the analysis of 54 responses to a survey and 312 community discussions shows that, while some users find badges somewhat appealing and nice to display, most find them unreliable as visual cues of developers’ activity or skills.
@@ 2.1 Social Translucence
One of the reasons behind the success of GitHub is that it has been designed as a ‘socially translucent’ system (Erickson and Kellogg, 2000), which makes socially salient information (e.g., stars, followers) and activities of participants (e.g., contribution calendar) visible, thus providing the basis for inferences, planning, and coordination. On the contrary, previous OSS forges like SourceForge were generally opaque to social information.
@@
We argue that profile pages, which show developers’ salient personal information, activity traces, and now also personal badges, act as a hub for socially translucent111Erickson and Kellogg use the adjective translucent instead of transparent to highlight privacy concerns and constant tension between showing just enough and too much information on socially augmented platforms. We use the term consistently. signals that support visibility, awareness, and accountability in GitHub.
@@ 2.2 Signaling Theory
When choosing with whom to collaborate, especially online, much of what we want to know about others is not directly observable. Instead, we rely on signals, i.e., perceivable features and actions that indicate the presence of the qualities of interest (Donath, 2007). For example, on collaborative development platforms like GitHub we can rely on the overall number of code-review comments as a signal of a developer’s timeliness and commitment; similarly, a large number of contributed pull requests can signal technical skills and high productivity.
@@
Signaling theory, initially developed in the fields of economics (Spence, 1978) and biology (Zahavi, 1975), is useful to model the relationship between signals and qualities, and to explain how only certain signals can be considered reliable. At its core, signaling theory is fundamentally concerned with reducing information asymmetry between the signaler—i.e., the party who possesses some pieces of information/personal attributes and benefits from the actions taken by the other party—and the receiver—i.e., an outsider who lacks information and would in turn benefit from making decisions based on it. For example, in biology, a peacock (the signaler) shows off the size and shape of his tail to be selected in favor of alternative partners (the signaler’s benefit) by a peahen (the receiver) because a larger tail is a signal of a healthy bird and a better chance of healthy offspring (the receiver’s benefit) (Zahavi, 1975).
@@
However, not all signals are reliable. To be considered reliable, the cost of deceiving a signal must outweigh the potential benefits. In terms of reliability, we distinguish two classes of signals. The assessment signals, which are inherently reliable because the signal can be produced only if the signaler possesses the indicated quality. For example, being able to lift a heavy weight in a gym is an assessment signal of someone’s strength—a person without enough strength simply would not be able. The other class of signals, named conventional signals, is not inherently reliable because the link between signal and quality is arbitrary, and is established by social conventions. For example, owning a gym t-shirt is a conventional signal of strength, as it is easily accessible, even if the wearer is weak. These types of signal are widespread in online interaction; for example, the descriptions on online profile pages are conventional signals, which are unreliable because they are more open to deception. Therefore, for signals to be reliable, they have to be observable—receivers need to be able to notice them— and they have to be costly to produce. In the labor market scenario, where employers lack information on job applicants, high-quality prospective employees distinguish themselves from the other low-quality applicants by highlighting in their resumé their education; holding a degree with honors from a prestigious institution is a signal that is easily observable for recruiters and is also costly to produce and hard to fake without being smart and hard working (Spence, 1978).
@@
Previous research has shown that easily observable signals conveyed through online profiles in translucent, social coding platforms (e.g., Stack Overflow reputation points, GitHub activity graph) can act as proxies of expertise (Shami et al., 2009; Marlow et al., 2013), awareness (Dabbish et al., 2012), and commitment (Tsay et al., 2013, 2014). Accordingly, here we want to investigate the extent to which personal badges in GitHub can be considered reliable as signals (whether assessment or conventional) of desirable yet directly unobservable qualities and personal dispositions of developers (e.g., timeliness, popularity, desire to collaborate, etc.), useful in the scenarios of open source software development collaboration (i.e., ‘people sensemaking’) (Shami et al., 2009) and developer hiring (Marlow and Dabbish, 2013).
@@ 2.3 Gamification in SE
Gamification in software engineering is common, especially in the social programmer software ecosystem (Singer et al., 2013), where several types of elements of game design are often used (de Paula Porto et al., 2021). Pedreira et al. (2015) conducted a systematic mapping study on gamification in software engineering and found that badges, i.e., tokens that users can display only after completing specified activities (Anderson et al., 2013; Fanfarelli et al., 2015), is the second most used gamification element after point-based reputation. For example, Stack Overflow implements a reward system that combines reputation scores and leaderboards with badges to publicly reward users for their contributions to the Q&A community.
GitHub has recently implemented personal badges to reward developers for the achievements unlocked through their activities on the coding platform. This is the first work to study the effects of badges and the reactions of the community to this new gamification element since its implementation in GitHub. Previous research on GitHub, in fact, has focused on badges as visual cues added by developers to README pages to give a quick look at repository and source code statistics (Trockman et al., 2018; Legay et al., 2019; Venigalla et al., 2022).
Although intended to increase user participation and steer user behavior in desired directions, gamification can have unwanted effects, even resulting in lower engagement, satisfaction, and performance (Hammedi et al., 2021). Moldon et al. (2021) examined how the behavior of GitHub developers changed after the removal of the code streak counters from the activity graphs displayed on their profile pages. They found that long-running streaks became less common, thus highlighting the power of gamification as a channel for social influence and raising awareness on the potential consequences of adding gamification elements to collaborative software development platforms.
Finally, Stack Overflow provides a complete list of existing achievements, as well as the rules and even community guidelines (Calefato et al., 2018) for detailing how to obtain them. However, an undesirable side effect of knowing which achievements are there and how to unlock them is that the behavior of users may change after obtaining a badge. Grant and Betts (2013) observed that the number of edits to posts in Stack Overflow drastically drops once users obtain related achievements. In GitHub, neither the list of achievements nor the rules for obtaining them are publicly disclosed. We speculate that this design choice may be intended to help avoid the risk that making the list of achievements public steers the developers’ coding behavior toward the game and, conversely, ensures that the regular coding activity on the platform is rewarded with achievement badges.
@@ 3 Research Questions
The overall goal of this paper is to characterize the diffusion, usefulness, and perception of personal badges in GitHub. Consequently, we define and answer the following four research questions.
RQ – What is the distribution of personal badges unlocked on GitHub?
With the first research question, we want not only to uncover the distribution of unlocked personal badges by type but also to assess whether and how many developers opted out of displaying them on their public profile page.
@@
The second research question builds on the signaling theory.
Because there is no official description of personal badges, we first seek to understand what developers’ qualities they are signaling (e.g., timeliness, popularity, desire to collaborate).
@@
RQ – Did the introduction of personal badges affect developers’ activities?
The unlocking rules and the badge list are not disclosed by GitHub.
However, there are several unofficial resources, such as repositories222https://github.com/Schweinepriester/github-profile-achievements and community discussions,333e.g., https://github.com/community/community/discussions/28656 which provide insights on the badges discovered so far and how to obtain them.
As such, we seek to understand the effects of the introduction of personal badges on the activities that developers typically perform in GitHub.
@@
Finally, with the fourth research question, we investigate the response of the GitHub community to the introduction and perceived usefulness of personal badges.
@@ 4.1 Data
Data were collected from a sample of organizations used in the preparation of the follow-up to a previous study of core developers who abandon GitHub projects (Calefato et al., 2022).
@@
Consequently, we started with the selection of organizations and, from the Topics section on the GitHub website, we identified the ten most trending topics.
Then, we moved on to select five to six organizations per topic.
We ended up selecting 58 organizations.
Finally, for each organization, we used the GitHub API to retrieve the list of organization members.
The API returned an error when retrieving the members of 7 organizations; hence, we ended up with 6,022 developers from 51 organizations.
Then, for each organization, we selected either ‘reference project’ (e.g.,
rails/rails,
nodejs/node); when unclear, we chose the largest in terms of contributors (i.e.,
laravell/framework) and, if more projects had a similar number of contributors, we chose the one with most stars.
@@
Refer to Table 5 in A for a breakdown of the characteristics of the selected projects.
Next, we detail the data collected and measures for each research question. To investigate RQ (diffusion of personal badges), because the GitHub API does not support the retrieval of personal badges, we developed a custom scraper in Python. For each user, we scraped the profile page to retrieve the list of unlocked personal badges and the time when the achievement was first obtained. GitHub introduced personal badges in June 2022; we repeated scraping monthly for six months, from June to December. We also counted how many GitHub developers chose not to have a public profile page or opted out of displaying personal badges from their public profile page.
To investigate the reliability of badges as signals of developers’ qualities (RQ) and the potential effects of badges on developers’ activities (RQ), we collected several repeated measures (see Section 4.2) concerning their development activities and popularity. Regarding development-related activities, we used the GitHub API to mine the number of Issues and Pull Requests (PRs) opened, closed (merged), and worked on (as assignee) by developers in our sample, the time to close issues and merge pull requests, and the number of commits. As a measure of the popularity of developers in the community GitHub, we also collected the number of followers. All these metrics were collected monthly for twelve months, from January to December 2022.
To gauge the perceptions of personal badges in GitHub (RQ), we used Google Forms to design an online survey consisting of 15 questions (both closed- and open-ended).
Survey questions (reported in B) are partially based on similar work by Trockman et al. (2018) who investigated repository badges in GitHub.
@@
The survey was advertised on social media and posted as a question to the GitHub community discussions space.444https://github.com/community/community/discussions/37346
@@
An anonymized copy of the survey is available in the replication package.555https://doi.org/10.5281/zenodo.7501582
Furthermore, to complement the survey and also further assess how the GitHub community reacted to the introduction of personal badges, we mined the questions and answers posted to the discussions space of the
community/community repository.
Specifically, we retrieved all posts listed in the profile category666https://github.com/community/community/discussions/categories/profile that also contained the keywords badge or achievement.

### 02 Feb 2024 v3

@@ Abstract
Context. GitHub has introduced a new gamification element through personal achievements, whereby badges are unlocked and displayed on developers’ personal profile pages in recognition of their development activities. Objective. In this paper, we present an exploratory analysis using mixed methods to study the diffusion of personal badges in GitHub, in addition to the effects and reactions to their introduction. Method. First, we conduct an observational study by mining longitudinal data from more than 6,000 developers and performed correlation and regression analysis. Then, we conduct a survey and analyze over 300 GitHub community discussions on the topic of personal badges to gauge how the community responded to the introduction of the new feature. Results. We find that most of the developers sampled own at least a badge, but we also observe an increasing number of users who choose to keep their profile private and opt out of displaying badges. Additionally, badges are generally poorly correlated with developers’ qualities and dispositions such as timeliness and desire to collaborate. We also find that, except for the Starstruck badge (reflecting the number of followers), their introduction does not have an effect. Finally, the reaction of the community has been in general mixed, as developers find them appealing in principle but without a clear purpose and hardly reflecting their abilities in the current form. Conclusions. We provide recommendations to the designers of the GitHub platform on how to improve the current implementation of personal badges as both a gamification mechanism and as sources of reliable cues for the assessment of developers’ abilities.
@@ 1 Introduction
Gamification refers to the practice of incorporating elements and mechanics of game design into non-game contexts (Hamari, 2007; Hunter and Werbach, 2012). The goal of gamification is to increase user engagement and motivation, as well as drive desired behaviors by creating a sense of play and/or competition (Deterding et al., 2011; Robson et al., 2015). There have been several attempts to propose frameworks for the gamification of Software Engineering activities (García et al., 2017; Sasso et al., 2017) and methods to build gamified software (Morschheuser et al., 2018). Previous research has also found evidence that gamification mechanisms are in part responsible for the success of websites such as Stack Overflow (SO) (Vasilescu et al., 2013). Gamification, in fact, has always been a staple of the technical Q&A site whose members gain badges, medals, and privileges by being good community citizens (e.g., answering questions, editing poorly written posts, etc.). Another popular source of user-generated content in the Software Engineering domain is GitHub, which, in contrast, has always incorporated several social features (e.g., developers have their personal profiles and can follow each other, repositories can be starred), but only implemented few and simple gamification elements on its website, such as code streak counters, which were removed from the contribution graph in May 2016 (Vessels, 2016). However, this suddenly changed in June 2022, when a blog post (Brooks, 2022) announced the introduction of personal achievements, a new feature whereby badges are unlocked in recognition of developers’ activities and displayed on their personal profile pages. However, unlike SO, neither the list of available badges is publicly disclosed by GitHub nor their meaning and the rules to unlock them. Previous research has shown that implementing gamification elements in a collaborative software platform may steer the behavior of developers in unexpected and unwanted directions (Moldon et al., 2021). We have found evidence that it is not entirely clear to the GitHub community how to interpret the meaning of personal achievements. In Figure 1, a user asks the community for help (1) on how to interpret the Pull Shark badge (2), and two other users agree with the original poster that the meaning of the badge is ambiguous and not intuitive (3).
@@
GitHub not only has earned the reputation of being the one-stop-shop for software development (Metz, 2015), but has also quickly become a useful platform for job seekers (sup, 2022; hac, 2011; ant, 2012). Employers access the GitHub profile pages of job candidates to supplement resumes and further assess their knowledge, skills, and abilities (Marlow and Dabbish, 2013). Marlow and Dabbish (2013) found that the design of GitHub strongly influenced the activity traces available through profile pages, which employers look at during applicant evaluation. In particular, they observed that employers look at the activity traces that provide reliable cues and require less effort to access and evaluate them. Because personal achievement badges are gamification elements that provide quick and low-effort access to activity traces, GitHub users who display them on their profile pages may be unintentionally ‘giving off’ cues that have unexpected repercussions, for example, to managers assessing them as job candidates and to other OSS developers who review their contributions (Singer et al., 2013; Tsay et al., 2014).
@@
In this paper, we investigate the effects of introducing a new gamification element into GitHub. Specifically, we conduct a mixed-methods study of personal achievement badges (in short, personal badges hereafter). We characterize the distribution of badges in GitHub (RQ), investigate the qualities of developers they are intended to signal and how well they correlate with these qualities (RQ), and examine the effects, if any, that the introduction of this feature has had on developers’ activities (RQ). Finally, we gauge how the GitHub community feels about the new feature (RQ).
@@
After mining longitudinal data from over 6,000 GitHub developers, we find that while most of them own at least one badge, the number of those who chose to have a private page or opted out of displaying badges on their profile has been steadily increasing. Furthermore, correlation and regression analyses reveal that, respectively, only one out of the five types of personal badges analyzed (i.e., Starstruck) provides a reliable signal of the associated developer quality (i.e., popularity) and that its introduction has positively affected the number of followers for developers displaying the badge. Finally, the analysis of 54 responses to a survey and 312 community discussions shows that, while some users find badges somewhat appealing and nice to display, most find them unreliable as visual cues of developers’ activity or skills.
@@ 2.1 Social Translucence
One of the reasons behind the success of GitHub is that it has been designed as a ‘socially translucent’ system (Erickson and Kellogg, 2000), which makes socially salient information (e.g., stars, followers) and the activities of participants (e.g., contribution calendar) visible, thus providing the basis for inferences, planning, and coordination. On the contrary, previous OSS forges like SourceForge were generally opaque to social information.
@@
We argue that profile pages, which show developers’ salient personal information, activity traces, and now also personal badges, act as a hub for socially translucent111Erickson and Kellogg use the adjective translucent instead of transparent to highlight privacy concerns and the constant tension between showing just enough and too much information on socially augmented platforms. We use the term consistently. signals that support visibility, awareness, and accountability in GitHub.
@@ 2.2 Signaling Theory
When choosing with whom to collaborate, especially online, much of what we want to know about others is not directly observable. Instead, we rely on signals, i.e., perceivable features and actions that indicate the presence of the qualities of interest (Donath, 2007). For example, on collaborative development platforms like GitHub we can rely on the total number of code-review comments as a signal of a developer’s timeliness and commitment; similarly, a large number of contributed pull requests can signal technical skills and high productivity.
@@
Signaling theory, initially developed in the fields of economics (Spence, 1978) and biology (Zahavi, 1975), is useful to model the relationship between signals and qualities, and to explain how only certain signals can be considered reliable. At its core, signaling theory is fundamentally concerned with reducing information asymmetry between the signaler – i.e., the party who possesses some pieces of information/personal attributes and benefits from the actions taken by the other party – and the receiver – i.e., an outsider who lacks information and would in turn benefit from making decisions based on it. For example, in biology, a peacock (the signaler) shows off the size and shape of his tail to be selected in favor of alternative partners (the signaler’s benefit) by a peahen (the receiver) because a larger tail is a signal of a healthy bird and a better chance of healthy offspring (the receiver’s benefit) (Zahavi, 1975).
@@
However, not all signals are reliable. To be considered reliable, the cost of deceiving a signal must outweigh the potential benefits. In terms of reliability, we distinguish two classes of signals. Assessment signals, which are inherently reliable because the signal can be produced only if the signaler possesses the indicated quality. For example, being able to lift a heavy weight in a gym is an assessment signal of someone’s strength – a person without enough strength would simply not be able. The other class of signals, named conventional signals, is not inherently reliable because the link between signal and quality is arbitrary, and is established by social conventions. For example, owning a gym t-shirt is a conventional signal of strength, as it is easily accessible, even if the wearer is weak. These types of signal are widespread in online interaction; for example, the descriptions on online profile pages are conventional signals, which are unreliable because they are more open to deception. Therefore, for signals to be reliable, they have to be observable – receivers need to be able to notice them – and they have to be costly to produce. In the labor market scenario, where employers lack information on job applicants, high-quality prospective employees distinguish themselves from the other low-quality applicants by highlighting in their resumé their education; holding a degree with honors from a prestigious institution is a signal that is easily observable for recruiters and is also costly to produce and hard to fake without being smart and hard working (Spence, 1978).
@@
Previous research has shown that easily observable signals conveyed through online profiles in translucent, social coding platforms (e.g., SO reputation points, GitHub activity graph) can act as proxies of expertise (Shami et al., 2009; Marlow et al., 2013), awareness (Dabbish et al., 2012), and commitment (Tsay et al., 2013, 2014). Consequently, here we investigate the extent to which personal badges in GitHub can be considered reliable as signals (whether assessment or conventional) of desirable yet directly unobservable qualities and personal dispositions of developers (e.g., timeliness, popularity, desire to collaborate, etc.), useful in the scenarios of open source software development collaboration (i.e., ‘people sensemaking’) (Shami et al., 2009) and developer hiring (Marlow and Dabbish, 2013).
@@ 2.3 Gamification in SE
Gamification in software engineering is common, especially in the social programmer software ecosystem (Singer et al., 2013), where several types of elements of game design are often used (de Paula Porto et al., 2021). Pedreira et al. (2015) conducted a systematic mapping study on gamification in software engineering and found that badges, i.e., tokens that users can display only after completing specified activities (Anderson et al., 2013; Fanfarelli et al., 2015), are the second most used gamification element after point-based reputation. For example, SO implements a reward system that combines reputation scores and leaderboards with badges to publicly reward users for their contributions.
GitHub has recently implemented personal badges to reward developers for achievements unlocked through their activities on the coding platform. This is the first work to study the effects of badges and the reactions of the community to this new gamification element since its implementation in GitHub. Previous research on GitHub, in fact, has focused on badges added by developers to README pages as visual cues that provide quick access to repository and source code statistics (Trockman et al., 2018; Legay et al., 2019; Venigalla et al., 2022).
Although intended to increase user participation and steer user behavior in desired directions, gamification can have unwanted effects, even resulting in lower engagement, satisfaction, and performance (Hammedi et al., 2021). Moldon et al. (2021) examined how the behavior of GitHub developers changed after the removal of the code streak counters from the activity graphs displayed on their profile pages. They found that long-lasting streaks became less common, thus highlighting the power of gamification as a channel of social influence and raising awareness on the potential consequences of adding gamification elements to collaborative development platforms.
Finally, SO provides a complete list of existing achievements, as well as rules and even community guidelines (Calefato et al., 2018) to detail how to obtain them. However, an undesirable side effect of knowing which achievements are there and how to unlock them is that the behavior of users may change after obtaining a badge. Grant and Betts (2013) observed that the number of edits to posts in SO drastically drops once users obtain related achievements. On GitHub, neither the list of achievements nor the rules for obtaining them are publicly disclosed. We speculate that the choice of not making the list of achievements public may be intended to help avoid the risk of steering developers’ behavior toward the game, thus ensuring that the regular coding activity on the platform is rewarded with achievement badges.
@@ 3 Research Questions
The overall goal of this paper is to characterize the diffusion, effects, and perception of personal badges in GitHub. Consequently, we define and answer the following four research questions.
RQ – What is the distribution of personal badges on GitHub?
With the first research question, we want not only to uncover the distribution of unlocked personal badges by type, but also to assess whether and how many developers opted out of displaying them on their public profile page.
@@
The second research question is based on signaling theory.
Because there is no official description of personal badges and how to obtain them, only unofficial resources such as repositories222https://github.com/Schweinepriester/github-profile-achievements and community discussions,333https://github.com/community/community/discussions/28656 we first seek to understand what developer qualities they are signaling (e.g., timeliness, popularity, desire to collaborate).
@@
RQ – Has the introduction of personal badges affected developers’ activities?
Although the ultimate benefit of implementing gamification in GitHub should be the promotion of best practices for collaborative software development, previous research has shown that gamification elements can instead lead to unintended changes in developers’ behavior.
As such, our objective is to understand the effects of the introduction of personal badges on the activities that developers typically perform in GitHub.
@@
Finally, we investigate the response of the GitHub community to the introduction and perceived usefulness of personal badges.
@@ 4.1 Data
Data were collected from a sample of organizations used in the preparation of the follow-up study on core developers who abandon GitHub projects (Calefato et al., 2022).
@@
Consequently, we started with the selection of organizations and,
First, from the Topics section on the GitHub website, we identified the ten most trending topics.
Five to six organizations per topic were selected, resulting in a total of 58 organizations.
For each them, we used the GitHub API to retrieve the list of organization members.
The API returned an error when retrieving the members of seven organizations; hence, we ended up with 6,022 developers from 51 organizations.
Then, for each organization, we selected the ‘reference project’ (e.g.,
rails/rails); when unclear (e.g.,
laravell/framework), we chose the largest in terms of contributors and, if more projects had a similar number of contributors, we chose the one with most stars.
@@
Refer to the supplementary material in the replication package444https://doi.org/10.5281/zenodo.7501582 for a breakdown of the selected projects’ characteristics.
Next, we detail the data collected and measures for each research question. To investigate RQ (diffusion of personal badges), because the GitHub API does not support the retrieval of personal badges, we developed a custom scraper in Python. For each user, we scraped the profile page to retrieve the list of unlocked personal badges and the time the achievement was first obtained. GitHub introduced personal badges in June 2022; we repeated scraping monthly for six months, from June to December. We also counted how many GitHub developers chose not to have a public profile page or opted out of displaying personal badges from their public profile page.
To investigate the reliability of badges as signals of developers’ qualities (RQ) and the potential effects of badges on developers’ activities (RQ), we collected several repeated measures (see Section 4.2) concerning their development activities and popularity. Regarding development-related activities, we used the GitHub API to mine the number of Issues and Pull Requests (PRs) opened, closed (merged), and worked on (as assignee) by developers in our sample, the time to close issues and merge pull requests, and the number of commits. As a measure of the popularity of developers in the GitHub community, we also collected the number of followers. All these metrics were collected monthly for twelve months, from January to December 2022.
To gauge the perceptions of personal badges in GitHub (RQ), we used Google Forms to design an online survey consisting of 15 questions (both closed and open).
@@
The survey was advertised on social media and posted as a question to the GitHub community discussions space.555https://github.com/community/community/discussions/37346
@@
A copy of the survey is available as supplementary material in the replication package.
Furthermore, to complement the survey and also further assess how the GitHub community reacted to the introduction of personal badges, we extracted all the discussions from the
community/community repository that were listed under the profile category666https://github.com/community/community/discussions/categories/profile and also contained the keyword badge and/or achievement.